OpenStudy (anonymous):

what is the solution of this equation? -8y+4= -20. check my answer

OpenStudy (anonymous):

do i subtract 4 from -20 and 4 or 4 and -8?

Nnesha (nnesha):

yes subtract 4 both side then divide by -8

OpenStudy (anonymous):

ohh okay hold on

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

Yep :)

Nnesha (nnesha):

nope a little mistake

OpenStudy (anonymous):

yeah it said wrong lol

OpenStudy (anonymous):

Ooh wait. I see CX

Nnesha (nnesha):

you subtract -20-4 what you get

OpenStudy (anonymous):

-24

Nnesha (nnesha):

you should get -24 so when you divide negative number by negative what your answer should be

OpenStudy (anonymous):

positive lol

OpenStudy (anonymous):

Negative divided by negative=postitve :)

Nnesha (nnesha):

yes so -24/-8 is positive 3

OpenStudy (anonymous):

Yep^^^

OpenStudy (anonymous):

2z=16 uhm what times 2 is 16??

OpenStudy (anonymous):

i divided and got 0.125

Nnesha (nnesha):

you have do divide 2 bot hside

OpenStudy (anonymous):

2 and 8 lol so then 4??

OpenStudy (anonymous):

well 16 divided by 2 is 8 and 2 divided by 8? is 4

Nnesha (nnesha):

no you have to divide only 2 both side then you done

OpenStudy (anonymous):

yeah 16/2 = 8

OpenStudy (anonymous):

so its 8?

Nnesha (nnesha):

yes right and left side 2/2 cancel out so z = 8 right

OpenStudy (anonymous):

lol Nnesha those medals

Nnesha (nnesha):

i just saw lol

OpenStudy (anonymous):

what about 1/2 y=1? im not sure if this one is stupid or not and haha yeah same